What is the amount of money that Lebo has to invest, at a simple interest rate of 12% per year, to earn R5250,00 interest after three years?


1
Expert's answer
2021-10-06T14:27:53-0400

I"=" P"\times" R"\times" T

5250"=" x"\times" 0.12"\times" 3

5250"=" 0.36x

X"=" "\frac{5250}{0.36}"

X"=" 14588,33

Answer is 14588,33